OpenStudy (anonymous):

i have a question. okay so the thing is 9^2/9^7. I know that i leave the number and subtract 7-2=5. so it would be 9^5. but do i but it as a fraction like 1/9^5 or do i just leave it

OpenStudy (anonymous):

put*

OpenStudy (anonymous):

second one

OpenStudy (popitree):

9^(-5)

OpenStudy (anonymous):

i just leave it?

OpenStudy (anonymous):

\[\frac{9^2}{9^7}=\frac{1}{9^{7-2}}=\frac{1}{9^5}\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw